OpenStudy (anonymous):

tan(0)=15/8 What does this mean?

OpenStudy (anonymous):

i think it must mean that \[\tan(\theta)=\frac{15}{8}\] because \[\tan(0)=0\]

OpenStudy (atlas):

Yeah right

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture where \[\tan(\theta)=\frac{15}{8}\] if you want the hypotenuse of the triangles solve \[h=\sqrt{15^2+8^2}=\sqrt{225+64}=\sqrt{289}=17\]

OpenStudy (anonymous):